Ask your own question, for FREE!
Mathematics 16 Online
OpenStudy (anonymous):

A bag contains 2 blue and 3 green marbles. Two marbles are drawn at random with replacement. Which of the following gives the probability of drawing two blue marbles?

OpenStudy (nikolas):

Try drawing out a tree diagram for these kinds of problems if it helps. For the first draw there are 2 blue marbles, and 5 marbles in total, so the probability that you will draw a blue marble is 2/5. For the next draw, if you already removed one blue marble, there is 1 blue marble remaining and 4 blue marbles in total, so the probability that you draw another blue marble is 1/4. You can multiply these two probabilities together to get the total probability of drawing two: 1/5 * 2/4 = 2/20 = 1/10. This exact problem is described, along with more helpful info, here: http://www.mathsisfun.com/data/probability-events-conditional.html

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!